A straight wire, 30.0 m long, moves at 2.0 m/s perpendicularly through a 1.0 T magnetic
field.
a. What EMF is induced in the wire?
b. The total resistance of the circuit of which the wire is a part is 15.0 Ω. What is the
current?
1
Expert's answer
2011-08-04T11:14:28-0400
a. Due to the Faraday's law, & EMF = vBL = 2.0 [m/s] 30.0 [m] 1.0 [T] = 60 [V]
b. Due to the Ohm's law: I = U/R = 60 [V]/15.0 [Ω] = 2 A.